# Spike-frequency and h-current based adaptation are dynamically equivalent in a Wilson-Cowan field model

### v2

@@ II.3 Wavenumbers
If a fixed point is identified as Turing- or Hopf-unstable, we compute the smallest wavenumber $k_{0}>0$ at which one of the conditions listed in Table 2 is met. Examples are shown in Figure 1, where $k_{0}$ is marked by the red circles. Additionally, we determine the wavenumber $k_{max}$ (see pink circles in Figure 1) for which the corresponding unstable mode grows most rapidly, i.e., for which the real part $\alpha_{j}(k)$ of the largest eigenvalue becomes maximal, $k_{max}=\text{argmax}_{k}(\max_{j}\alpha_{j}(k))$. Wavenumbers were computed, because they determine the wavelengths of the spatio-temporal patterns that emerge in a fixed point that undergoes one of the aforementioned bifurcations (see [23, 63, 43]).
@@ II.4 Symmetries and equivalence of mechanisms
be the system with SFA (i.e., let $b,\ \beta_{m}>0$), and let $T:[0,1]^{3}\to[0,1]^{3}$ with $T:(u_{e},u_{i},m)^{\top}\mapsto(u_{e},u_{i},1-m)^{\top}$ be an isomorphic transformation. Then we have $\frac{\partial}{\partial t}\bigl[T\bigl((u_{e}(x,t),u_{i}(x,t),m(x,t))^{\top}\bigr)\bigr]=(\frac{\partial u_{e}}{\partial t},\frac{\partial u_{i}}{\partial t},-\frac{\partial m}{\partial t})^{\top}$. Inserting the transformation into Equations (10) we obtain
$S(\mathbf{T(u_{e},u_{i},m)})=\begin{pmatrix}\tau_{e}\frac{\partial u_{e}}{\partial t}\\ \tau_{i}\frac{\partial u_{i}}{\partial t}\\ \mathbf{-\tau_{m}\frac{\partial m}{\partial t}}\end{pmatrix}$= $\begin{pmatrix}-u_{e}(x,t)+F_{e}(w_{ee}w_{e}(x)\ast u_{e}(x,t)-w_{ei}w_{i}(x)\ast u_{i}(x,t)-\mathbf{b(1-m(x,t))}+I_{e})\\ -u_{i}(x,t)+F_{i}(w_{ie}w_{e}(x)\ast u_{e}(x,t)-w_{ii}w_{i}(x)\ast u_{i}(x,t)+I_{i})\\ -\bigl(\mathbf{-(1-m(x,t)})+\underbrace{F_{m}(u_{e}(x,t)-\mu)}_{=1-F_{m}(\mu-u_{e})}\bigr)\end{pmatrix}$ (11)
$=$ $\begin{pmatrix}-u_{e}(x,t)+F_{e}(w_{ee}w_{e}(x)\ast u_{e}(x,t)-w_{ei}w_{i}(x)\ast u_{i}(x,t)+bm(x,t)-b+I_{e})\\ -u_{i}(x,t)+F_{i}(w_{ie}w_{e}(x)\ast u_{e}(x,t)-w_{ii}w_{i}(x)\ast u_{i}(x,t)+I_{i})\\ -m(x,t)+F_{m}(\mu-u_{e}(x,t))\end{pmatrix}$. (12)
This system of equations is, given the choice of $F_{m}$ (Equation (4)), equivalent to the system of Equations (1) for h-currents if $b\leftarrow b$ and $\beta_{m}\leftarrow\beta_{m}$ (see Table 1) and the external input to the excitatory population in the system with SFA is decreased by the value of $b$, i.e., $I_{e}^{h}=I_{e}^{a}-b$. The superscripts $\cdot^{a}$, $\cdot^{h}$ indicate the difference between the variables and parameters for SFA and h-currents, respectively. Hence, both adaptation mechanisms are dynamically equivalent under the transformation $T$, if $I_{e}$ is correspondingly changed and if $F_{m}^{a}(x)=1-F_{m}^{h}(x)$ holds for the adaptation nonlinearity. This equivalence also tells us that a fixed point $(\tilde{u}_{e}^{a},\tilde{u}_{i}^{a},\tilde{m}^{a})$ for the system with SFA corresponds to a fixed point $(\tilde{u}_{e}^{h},\tilde{u}_{i}^{h},1-\tilde{m}^{h})$ for the system with h-currents and that the dynamical landscape for h-currents should be the same as the dynamical landscape for SFA, but shifted along the $I_{e}$-axis by the value $b$.
Beyond dynamical equivalence between the adaptation mechanisms, we also have a symmetry for each system within itself. Let $T:[0,1]^{3}\to[0,1]^{3}$ with $T:(u_{e},u_{i},m)^{\top}\mapsto(1-u_{e},1-u_{i},1-m)^{\top}$ be another isomorphic transformation such that $\frac{\partial}{\partial t}\bigl[T\bigl((u_{e}(x,t),u_{i}(x,t),m(x,t))^{\top}\bigr)\bigr]=(-\frac{\partial u_{e}}{\partial t},-\frac{\partial u_{i}}{\partial t},-\frac{\partial m}{\partial t})^{\top}$. Inserting this in Equation (10) we obtain
@@
If $\hat{I}_{e}=-I_{e}+b-w_{ee}+w_{ei},\ \hat{I}_{i}=-I_{i}-w_{ie}+w_{ii}$, and $\hat{\mu}=1-\mu$, then the system (13) is equivalent to Equations (1) for both SFA and h-currents. This induces an approximate point symmetry of state space at $(I^{c}_{e},I^{c}_{i})=\left(\frac{b-w_{ee}+w_{ei}}{2},\frac{-w_{ie}+w_{ii}}{2}\right)$ between $(u_{e},u_{i},m)$ and $(1-u_{e},1-u_{i},1-m)$ which becomes exact for $\mu=0.5$. Since $\mu=0.4$ in this study, an exact symmetry is only achieved for $b=0$ (Figure 2A, first column, where the black dot in the middle of the orange regime denotes the center of the point reflection).
@@ III Overview of state space and its changes with adaptation strength
Figure 2 shows the results of the analytical investigation for the system with h-currents with increasing adaptation strength (Figure 2A) and examples for the emerging spatio-temporal patterns (Figure 2B) for all regimes of Hopf or Turing instability. Without adaptation ($b=0$), we observe a Hopf-unstable regime (orange) in which both temporal (similar to Figure 2B, panel 3(){(*)}, for b=0.5b=-0.5) and spatial (see Figure 2B, panel 1) patterns emerge. With adaptation, the Hopf regime is also apparent but the activity patterns can be either temporal (Figure 2B, panel 3()(*)) or spatio-temporal (Figure 2B, panel 3). Close to the bifurcation lines of the Hopf-unstable regime, we observe the emergence of Turing instability in down-states (upper border; for an activity example see Figure 2B, panel 4) and in up-states (lower border; for activity example see Figure 2B, panel 6). Beyond the Turing instability in down- (up-)states, to the left (right) side of the state spaces in Figure 2A, there exists a regime, marked dark blue (dark red), where a single down- (up-) state fixed point is stable. We also see regimes of three fixed points including one stable up-state (dark beige) while the other two fixed points are unstable and vice versa with one stable down-state (light beige). Activity traces in these regimes are similar to the activities shown in Figure 2B, panels 5 and 7. Initializations close to zero always converge to the existing stable fixed point while initializations around the unstable states cause irregular traveling waves. Further, to both corner ends (upper right and lower left), we find a regime of bistability between up- and down-states (light blue) next to bistable regimes, where a stable down-state coexists with a Turing instability in the up-state (dark purple) and vice versa (dark green). Figure 2C, panel 5, shows that the activity converges to the activity caused by the Turing bifurcation in the down-state, if initialized close to the Turing-unstable down-state. Initialized close to zero, the activity converges to the stable up-state (see activity 5(){(*)}). The pattern emerging in the Turing-unstable down-state of panel 5 shows down-state waves traveling over an otherwise stable up-state activity, unlike the traveling waves shown in panels 4 and 6, where no other stable fixed point coexists. Furthermore, in some states, the activity patterns can become irregular (see Figure 2B, panel 7), such that they do not converge to a periodic, equidistant, or otherwise regular pattern. Such irregular patterns predominantly occur at the border of different regimes or in states where another attractor, e.g., a stable fixed point, coexists.
@@
The Hopf-unstable states further undergo Turing bifurcations beyond the Hopf bifurcation at $k=0$ (see, for example, Figure 1, top row, third panel), giving rise to spatial (for $b=0$, static Turing bifurcations) or spatio-temporal (for $b\neq 0$, dynamic Turing bifurcations) patterns. Initialized close to zero, states in the Hopf regime converge to fast, spatially homogeneous, temporal oscillations caused by the Hopf instability, while initializations close to the fixed points converge to spatial or spatio-temporal patterns. In this regime, the system thus experiences a second type of bistability that is not caused by the coexistence of multiple fixed points [^4] Note, however, that in the following the name “bistable regime” will refer to the light blue area in state space, where two stable fixed points coexist.. However, there are two small areas in the Hopf-unstable regime for $b=0$ where exceptions occur and where spatio-temporal patterns are present (Figure 2B, panel 0). These areas are located in the top left of the Hopf regime where a stable down-state, three fixed points with a stable down-state, a Hopf instability, and a Turing instability are getting close in parameter space, and in the bottom right corner, where the regimes of a stable up-state, a Hopf instability, three fixed points with a stable up-state, and a Turing instability meet. In these states, depending on the initialization, most of the ring converges toward the standing wave interspersed by a few traveling bumps (see Figure 2B, panel 0). When adaptation strength increases, the bistable regime (light blue) shrinks and vanishes for values $|b|>1$. The size of the Hopf-unstable regime (orange) and of the regime of three fixed points with a stable up-state (beige) increase, while the regime of three fixed points with a stable down-state decreases in size (light beige).
@@
Figure 3 shows a comparison between the state spaces and activity traces for the systems with SFA and h-currents. We see in Figure 3A that the dynamical regimes are equivalent for SFA (left) and h-currents (right), except for the shift to the right (left) by the absolute value $|b|$ for SFA (h-currents), consistent with the required compensation of external inputs to the excitatory population (see Section II.4). The rightward shift of the dynamical regimes for SFA compared to h-currents implies that the system requires more external input to the excitatory population to produce activity patterns beyond the spatially and temporally homogeneous steady states. We also see that equivalent positions (considering the required compensatory inputs) cause activity patterns to be equivalent under the transformation $T$, i.e., for $m\leftarrow 1-m$, as exemplified in Figure 3B.
@@
Figure 4: Eigenvalue spectrum for small values of $|b|$. The panels show the real, $\alpha_{0,1,2}(k)$, (solid lines) and imaginary parts, $\omega_{0,1,2}(k)$, (dashed lines) of the eigenvalues of the linearization matrix in Equation (7). Top row: Hopf-unstable state at $(I_{e},I_{i})=(-0.35,-1.225)$. Center row: Turing-unstable down-state at $(I_{e},I_{i})=(-0.35,-0.55)$. Bottom row: same Turing-unstable down-state as in the center row but with weaker adaptation. Adaptation strength $|b|$ decreases from left to right. Lower panels show close-ups around the zero crossing of the real parts. The imaginary parts vanish for $b=0$. Arrows in panels below the first two plots of the upper row indicate zero crossing of a static (s) and dynamic (d) Turing bifurcation.
@@
For the investigated parameterization (see Table 1), adaptation is required to cause the purely temporal or spatial patterns observed at $b=0$ to travel. We exemplify this in Figure 4 where we show the real ($\alpha_{j}(k)$) and imaginary ($\omega_{j}(k)$) parts of all three eigenvalues $\lambda_{j}(k)=\alpha_{j}(k)\pm i\omega_{j}(k)$ of the Jacobian (Equation (7)) for a Hopf-unstable (upper row) and a Turing-unstable (center and bottom row) state. In the upper two rows, for small values of $b$, the zero crossings of the real parts of any eigenvalue coincide with non-zero imaginary parts, resulting in a dynamic Turing bifurcation and enabling spatio-temporal patterns in that very state. For $b=0$ the imaginary part vanishes, resulting in a static Turing bifurcation. However, adaptation can not be arbitrarily weak and must be stronger than some threshold to show dynamic Turing bifurcations. In the bottom row of Figure 4, for example, the Turing bifurcations are static for b=0.02b=-0.02, and b=0.01b=-0.01 and allow the system to produce standing waves (see Appendix A.3, Figure 8). While the Turing bifurcation is dynamic for b=0.03b=-0.03, all eigenvalues have negative real parts for b=0.02b=-0.02 for every $k$ where the imaginary part is non-zero, i.e., the system is stable whenever there is at least one eigenvalue $j\in\{0,1,2\}$ such that $\omega_{j}(k)\neq 0$.
@@
Additionally, we see in Figure 4, as $|b|$ increases, the imaginary part becomes larger which indicates faster temporal oscillations, since each mode is given by
$\exp(\lambda(k)t)\ \exp(ikx)=\exp((\underbrace{\alpha(k)}_{=0}\pm i\omega(k))\ \exp(ikx)=\exp(\pm i\underbrace{\omega(k)}_{\mathclap{\omega\uparrow\text{ for }|b|\uparrow}}t)\ \exp(ikx).$
@@
Above conclusions are confirmed for the entire slices of state space in Appendix A.1, Figure 7, where we distinguish both types of bifurcations that occur.

### v3

@@ II.3 Wavenumbers
If a fixed point is identified as Turing- or Hopf-unstable, we compute the smallest wavenumber $k_{0}>0$ at which one of the conditions listed in Table 2 is met. Examples are shown in Fig. 1, where $k_{0}$ is marked by the red circles. Additionally, we determine the wavenumber $k_{max}$ (see pink circles in Fig. 1) for which the corresponding unstable mode grows most rapidly, i.e., for which the real part $\alpha_{j}(k)$ of the largest eigenvalue becomes maximal, $k_{max}=\text{argmax}_{k}(\max_{j}\alpha_{j}(k))$. Wavenumbers were computed, because they determine the wavelengths of the spatio-temporal patterns that emerge in a fixed point that undergoes one of the aforementioned bifurcations (see [23, 63, 43]).
@@ II.4 Symmetries and equivalence of mechanisms
be the system with SFA (i.e., let $b,\ \beta_{m}>0$), and let $T:[0,1]^{3}\to[0,1]^{3}$ with
$T:(u_{e},u_{i},m)^{\top}\mapsto(u_{e},u_{i},1-m)^{\top}$
be an isomorphic transformation. Then we have $\frac{\partial}{\partial t}\bigl[T\bigl((u_{e}(x,t),u_{i}(x,t),m(x,t))^{\top}\bigr)\bigr]=(\frac{\partial u_{e}}{\partial t},\frac{\partial u_{i}}{\partial t},-\frac{\partial m}{\partial t})^{\top}$. Inserting the transformation into Eqs. (8) we obtain
$S(\mathbf{T(u_{e},u_{i},m)})=\begin{pmatrix}\tau_{e}\frac{\partial u_{e}}{\partial t}\\ \tau_{i}\frac{\partial u_{i}}{\partial t}\\ \mathbf{-\tau_{m}\frac{\partial m}{\partial t}}\end{pmatrix}$= $\begin{pmatrix}-u_{e}(x,t)+F_{e}(w_{ee}w_{e}(x)\ast u_{e}(x,t)-w_{ei}w_{i}(x)\ast u_{i}(x,t)-\mathbf{b(1-m(x,t))}+I_{e})\\ -u_{i}(x,t)+F_{i}(w_{ie}w_{e}(x)\ast u_{e}(x,t)-w_{ii}w_{i}(x)\ast u_{i}(x,t)+I_{i})\\ -\bigl(\mathbf{-(1-m(x,t)})+\underbrace{F_{m}(u_{e}(x,t)-\mu)}_{=1-F_{m}(\mu-u_{e})}\bigr)\end{pmatrix}$
$=$ $\begin{pmatrix}-u_{e}(x,t)+F_{e}(w_{ee}w_{e}(x)\ast u_{e}(x,t)-w_{ei}w_{i}(x)\ast u_{i}(x,t)+bm(x,t)-b+I_{e})\\ -u_{i}(x,t)+F_{i}(w_{ie}w_{e}(x)\ast u_{e}(x,t)-w_{ii}w_{i}(x)\ast u_{i}(x,t)+I_{i})\\ -m(x,t)+F_{m}(\mu-u_{e}(x,t))\end{pmatrix}$.
This system of equations is, given the choice of $F_{m}$ (Eq. (3)), equivalent to the system of Eqs. (1) for h-currents if $b\leftarrow b$ and $\beta_{m}\leftarrow\beta_{m}$ (see Table 1) and the external input to the excitatory population in the system with SFA is decreased by the value of $b$, i.e., $I_{e}^{h}=I_{e}^{a}-b$. The superscripts $\cdot^{a}$, $\cdot^{h}$ indicate the difference between the variables and parameters for SFA and h-currents, respectively. Hence, both adaptation mechanisms are dynamically equivalent under the transformation $T$, if $I_{e}$ is correspondingly changed and if $F_{m}^{a}(x)=1-F_{m}^{h}(x)$ holds for the adaptation nonlinearity. This equivalence also tells us that a fixed point $(\tilde{u}_{e}^{a},\tilde{u}_{i}^{a},\tilde{m}^{a})$ for the system with SFA corresponds to a fixed point $(\tilde{u}_{e}^{h},\tilde{u}_{i}^{h},1-\tilde{m}^{h})$ for the system with h-currents and that the dynamical landscape for h-currents should be the same as the dynamical landscape for SFA, but shifted along the $I_{e}$-axis by the value $b$.
Beyond dynamical equivalence between the adaptation mechanisms, we also have a symmetry for each system within itself. Let $T:[0,1]^{3}\to[0,1]^{3}$ with $T:(u_{e},u_{i},m)^{\top}\mapsto(1-u_{e},1-u_{i},1-m)^{\top}$ be another isomorphic transformation such that $\frac{\partial}{\partial t}\bigl[T\bigl((u_{e}(x,t),u_{i}(x,t),m(x,t))^{\top}\bigr)\bigr]=(-\frac{\partial u_{e}}{\partial t},-\frac{\partial u_{i}}{\partial t},-\frac{\partial m}{\partial t})^{\top}$. Inserting this in Eqs. (8) we obtain
@@
If $\hat{I}_{e}=-I_{e}+b-w_{ee}+w_{ei},\ \hat{I}_{i}=-I_{i}-w_{ie}+w_{ii}$, and $\hat{\mu}=1-\mu$, then the system (9) is equivalent to Eqs. (1) for both SFA and h-currents. This induces an approximate point symmetry of state space at $(I^{c}_{e},I^{c}_{i})=\left(\frac{b-w_{ee}+w_{ei}}{2},\frac{-w_{ie}+w_{ii}}{2}\right)$ between $(u_{e},u_{i},m)$ and $(1-u_{e},1-u_{i},1-m)$ which becomes exact for $\mu=0.5$. Since $\mu=0.4$ in this study, an exact symmetry is only achieved for $b=0$ (Fig. 2A, first column, where the black dot in the middle of the orange regime denotes the center of the point reflection).
@@ III Overview of state space and its changes with adaptation strength
Figure 2 shows the results of the analytical investigation for the system with h-currents with increasing adaptation strength (Fig. 2A) and examples for the emerging spatio-temporal patterns (Fig. 2B) for all regimes of Hopf or Turing instability. Without adaptation ($b=0$), we observe a Hopf-unstable regime (orange) in which both temporal (similar to Fig. 2B, panel 3(){(*)}, for b=0.5b=-0.5) and spatial (see Fig. 2B, panel 1) patterns emerge. With adaptation, the Hopf regime is also apparent but the activity patterns can be either temporal (Fig. 2B, panel 3()(*)) or spatio-temporal (Fig. 2B, panel 3). Close to the bifurcation lines of the Hopf-unstable regime, we observe the emergence of Turing instability in down-states (upper border; for an activity example see Fig. 2B, panel 4) and in up-states (lower border; for activity example see Fig. 2B, panel 6). Beyond the Turing instability in down- (up-) states, to the left (right) side of the state spaces in Fig. 2A, there exists a regime, marked dark blue (dark red), where a single down- (up-) state fixed point is stable. We also see regimes of three fixed points including one stable up-state (dark beige) while the other two fixed points are unstable and vice versa with one stable down-state (light beige). Activity traces in these regimes are similar to the activities shown in Fig. 2B, panels 5(*) and 7. Initializations close to zero always converge to the existing stable fixed point while initializations around the unstable states cause irregular traveling waves. Further, to both corner ends (upper right and lower left), we find a regime of bistability between up- and down-states (light blue) next to bistable regimes, where a stable down-state coexists with a Turing instability in the up-state (dark purple) and vice versa (dark green). Figure 2C, panel 5, shows that the activity converges to the activity caused by the Turing bifurcation in the down-state, if initialized close to the Turing-unstable down-state. Initialized close to zero, the activity converges to the stable up-state (see activity 5(){(*)}). The pattern emerging in the Turing-unstable down-state of panel 5 shows down-state waves traveling over an otherwise stable up-state activity, unlike the traveling waves shown in panels 4 and 6, where no other stable fixed point coexists. Furthermore, in some states, the activity patterns can become irregular (see Fig. 2B, panel 7), such that they do not converge to a periodic, equidistant, or otherwise regular pattern. Such irregular patterns predominantly occur at the border of different regimes or in states where another attractor, e.g., a stable fixed point, coexists.
@@
The Hopf-unstable states further undergo Turing bifurcations beyond the Hopf bifurcation at $k=0$ (see, for example, Fig. 1, top row, third panel), giving rise to spatial (for $b=0$, static Turing bifurcations) or spatio-temporal (for $b\neq 0$, dynamic Turing bifurcations) patterns. Initialized close to zero, states in the Hopf regime converge to fast, spatially homogeneous, temporal oscillations caused by the Hopf instability, while initializations close to the fixed points converge to spatial or spatio-temporal patterns. In this regime, the system thus experiences a second type of bistability that is not caused by the coexistence of multiple fixed points [^5] Note, however, that in the following the name “bistable regime” will refer to the light blue area in state space, where two stable fixed points coexist.. However, there are two small areas in the Hopf-unstable regime for $b=0$ where exceptions occur and where spatio-temporal patterns are present (Fig. 2B, panel 0). These areas are located in the top left of the Hopf regime where a stable down-state, three fixed points with a stable down-state, a Hopf instability, and a Turing instability are getting close in parameter space, and in the bottom right corner, where the regimes of a stable up-state, a Hopf instability, three fixed points with a stable up-state, and a Turing instability meet. In these states, depending on the initialization, most of the ring converges toward the standing wave interspersed by a few traveling bumps (see Fig. 2B, panel 0). When adaptation strength increases, the bistable regime (light blue) shrinks and vanishes for values $|b|>1$. The size of the Hopf-unstable regime (orange) and of the regime of three fixed points with a stable up-state (beige) increase, while the regime of three fixed points with a stable down-state decreases in size (light beige).
@@
Figure 3 shows a comparison between the state spaces and activity traces for the systems with SFA and h-currents. We see in Fig. 3A that the dynamical regimes are equivalent for SFA (left) and h-currents (right), except for the shift to the right (left) by the absolute value $|b|$ for SFA (h-currents), consistent with the required compensation of external inputs to the excitatory population (see Section II.4). The rightward shift of the dynamical regimes for SFA compared to h-currents implies that the system requires more external input to the excitatory population to produce activity patterns beyond the spatially and temporally homogeneous steady states. We also see that equivalent positions (considering the required compensatory inputs) cause activity patterns to be equivalent under the transformation $T$, i.e., for $m\leftarrow 1-m$, as exemplified in Fig. 3B.
@@
Figure 4: Eigenvalue spectrum for small values of $|b|$. The panels show the real, $\alpha_{0,1,2}(k)$, (solid lines) and imaginary parts, $\omega_{0,1,2}(k)$, (dashed lines) of the eigenvalues of the linearization matrix in Eq. (6). Top row: Hopf-unstable state at $(I_{e},I_{i})=(-0.35,-1.225)$. Center row: Turing-unstable down-state at $(I_{e},I_{i})=(-0.35,-0.55)$. Bottom row: same Turing-unstable down-state as in the center row but with weaker adaptation. Adaptation strength $|b|$ decreases from left to right. Lower panels show close-ups around the zero crossing of the real parts. The imaginary parts vanish for $b=0$. Arrows in panels below the first two plots of the upper row indicate zero crossing of a static (s) and dynamic (d) Turing bifurcation.
@@
For the investigated parameterization (see Table 1), adaptation is required to cause the purely temporal or spatial patterns observed at $b=0$ to travel. We exemplify this in Fig. 4 where we show the real ($\alpha_{j}(k)$) and imaginary ($\omega_{j}(k)$) parts of all three eigenvalues $\lambda_{j}(k)=\alpha_{j}(k)\pm i\omega_{j}(k)$ of the Jacobian (Eq. (6)) for a Hopf-unstable (upper row) and a Turing-unstable (center and bottom row) state. In the upper two rows, for small values of $b$, the zero crossings of the real parts of any eigenvalue coincide with non-zero imaginary parts, resulting in a dynamic Turing bifurcation and enabling spatio-temporal patterns in that very state. For $b=0$ the imaginary part vanishes, resulting in a static Turing bifurcation. However, adaptation can not be arbitrarily weak and must be stronger than some threshold to show dynamic Turing bifurcations. In the bottom row of Fig. 4, for example, the Turing bifurcations are static for b=0.02b=-0.02 and b=0.01b=-0.01 and allow the system to produce standing waves for the corresponding activity traces (see Appendix A.3, Fig. 8). While the Turing bifurcation becomes dynamic for b=0.03b=-0.03, all eigenvalues have negative real parts for b=0.02b=-0.02 for every $k$ for which the imaginary part is non-zero, i.e., the system is stable whenever there is at least one eigenvalue $j\in\{0,1,2\}$ such that $\omega_{j}(k)\neq 0$.
@@
Additionally, we see in Fig. 4, as $|b|$ increases, the imaginary part becomes larger which indicates faster temporal oscillations, since each mode is given by
$\exp(\lambda(k)t)\ \exp(ikx)=$ $\ \exp((\underbrace{\alpha(k)}_{=0}\pm i\omega(k))\ \exp(ikx)$
$=$ $\ \exp(\pm i\underbrace{\omega(k)}_{\mathclap{\omega\uparrow\text{ for }|b|\uparrow}}t)\ \exp(ikx).$
@@
Above conclusions are confirmed for the entire slices of state space in Appendix A.1, Fig. 7, where we distinguish both types of bifurcations that occur.